JaiLai Cos. stock has a beta of 0.6, the current risk-free rate is 6.4 percent, and the expected return on the market is 12 percent.


What is JaiLai’s cost of equity? (Round your answer to 2 decimal places.)


  Cost of equity %
Answer #1

Using the CAPM model,

ERi = Rf + β(ERm - Rf)

where,

ERi = Expected cost of equity
Rf = Risk-free rate = 6.4%
β = Beta of the investment = 0.6
ERm = Expected return of market = 12%

=> ERi = 6.4 + 0.6(12 - 6.4) = 9.76%